I know that the number 243 is in exponential notation.
3^5 = 3򉁫򉁫 in exponential notation. I know this because there are calculators to tell me this. But HOW would I find this exponent without a calculator?
thank you so much. I'm here to understand.

Answer by scott8148(5896)

factor the number into prime factors __ prime numbers that can't be factored

243 is odd, so not divisible by 2

243/3=81 __ aha!! a perfect square __ so 243=3*9*9 __ 9's are factorable __ so 243=3*3*3*3*3